Denek Box company produces card boxes In bundles of 1,000 boxes. The market Is highly competitive, with boxes currently selling for $100 per thousand. Denek's total and marginal cost curves are:

TC = 3,000,000 + 0.001Q2

2

MC = 0.002Q

Q Is measured In thousand box bundles per year.

a. Calculate Denek's profit maximizing quantity.

b. Is the firm earning a profit/loss?

c. Analyze Denek's position in terms of the shutdown condition.


1
Expert's answer
2021-04-15T06:55:09-0400

a. For a perfectly competitive firm, profit maximizing quantity is given at the point where:

Price = Marginal Cost

100 = 0.002Q

"Q*= \frac{100}{0.002}=50000"

b. Profit = Total Revenue (TR) – Total Cost (TC)

"= PQ* - (3000000+0.001Q^2) \\

= 100 \times 50000 – 3000000 – 0.001(50000)^2 \\

= 5000000 – 3000000 – 2500000 \\

= -500000"

The firm is making a loss of $500000

c. A firm should operate in perfect competition as long as price is greater than average variable cost (AVC)

P>AVC

"AVC = \frac{Total \; variablr \; cost \;(TVC)}{Q} \\

= \frac{0.001Q^2}{Q} \\

= 0.001Q"

When Q* = 50000

AVC = "0.001 \times 50000 = 50"

Thus, P>AVC, so the firm should continue the production.